### v1

@@ Abstract
Remote sensing image (RSI) interpretation typically faces challenges due to the scarcity of labeled data, which limits the performance of RSI interpretation tasks. To tackle this challenge, we propose EarthSynth, a diffusion-based generative foundation model that enables synthesizing multi-category, cross-satellite labeled Earth observation for downstream RSI interpretation tasks. To the best of our knowledge, EarthSynth is the first to explore multi-task generation for remote sensing. EarthSynth, trained on the EarthSynth-180K dataset, employs the Counterfactual Composition training strategy to improve training data diversity and enhance category control. Furthermore, a rule-based method of R-Filter is proposed to filter more informative synthetic data for downstream tasks. We evaluate our EarthSynth on scene classification, object detection, and semantic segmentation in open-world scenarios, offering a practical solution for advancing RSI interpretation.
@@ 1 Introduction
In parallel, recent advances in generative data augmentation antoniou2017data (GDA), particularly diffusion models rombach2022high (DMs), offer a promising avenue for synthesizing high-quality labeled training data. GDA using DMs has been widely applied across various domains chen2024comprehensive . Prior works he2022synthetic ; trabucco2023effective investigate text-to-image (T2I) models to boost classification performance in data-scarce scenarios, enlightening and inspiring the remote sensing community. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. By augmenting rare classes and enriching data diversity, these models can play a critical role in mitigating data scarcity and enhancing the generalization capabilities of RSI interpretation.
@@
In the remote sensing community, most existing GDA approaches are trained predominantly on single-source EO, which inherently limits the diversity of generated categories and lacks generality and flexibility across broader remote sensing applications. However, training generative models on multi-source data requires balancing the quality and diversity of generated data. Txt2Img-MHN xu2023txt2img attempts to use GANs salimans2016improved to generate satellite images. DiffusionSat khanna2023diffusionsat and CRS-Diff tang2024crs propose conditional DMs for generating optical RSIs, incorporating diverse texture-based conditions and applying the synthesized data to downstream tasks such as road extraction. GeoSynth toker2024satsynth joint data manifold of images and labels for satellite semantic segmentation. AeroGen tang2024aerogen and MMO-IG yang2025mmo try to use synthesized training data for satellite object detection. However, these methods are typically trained on single tasks or single-source data, requiring repeated training and generation to meet diverse needs, which hinders their direct application in real-world scenarios.
@@
To solve the above problems, we propose EarthSynth, a diffusion-based generative foundation model, synthesizing EO with a semantic mask and text for downstream RSI interpretation tasks as shown in Figure 1. First, we construct the EarthSynth-180K with multi-source and multi-category data, 180K samples for training EarthSynth. Specifically, we collect samples from multiple datasets and apply random cropping and category-augmentation strategy to standardize image resolution, ensuring alignment among images, semantic masks, and text descriptions. To our knowledge, EarthSynth-180K is the first large-scale remote sensing dataset for diffusion training. During training, we adopt the Counterfactual Composition (CF-Comp) strategy to simultaneously enhance layout controllability and category diversity, thereby enabling the generation of more informative EO data. For training data synthesis, a rule-based method of R-Filter is proposed to filter more informative synthetic data. We evaluate our EarthSynth with multiple datasets on scene classification, object detection, and semantic segmentation. Furthermore, the effectiveness of our method is demonstrated through comprehensive ablation studies and visualization analysis.
@@
  • We propose EarthSynth, a diffusion-based generative foundation model trained on the EarthSynth-180K dataset with 180K samples aligned image, semantic mask, and text, achieving a unified solution to achieve multi-task generation.
  • EarthSynth employs the CF-Comp strategy to balance the layout controllability and category diversity during training, enabling fine layout control for RSI generation. It further incorporates the R-Filter to extract more informative and high-quality synthesized samples.
  • EarthSynth is evaluated on scene classification, object detection, and semantic segmentation, three RSI interpretation tasks in open-world scenarios.
@@ 2 Related Work
Diffusion Models for Remote Sensing. Diffusion models, which have shown great success in natural image synthesis rombach2022high , are increasingly being applied to remote sensing khanna2023diffusionsat ; liu2025text2earth . Depending on the use case liu2024diffusion , their applications in remote sensing can be broadly divided into three categories: image enhancement, image interpretation, and image synthesis. For image enhancement, DMs improve multispectral and hyperspectral images by fusing information across channels and expanding attention xu2023dual ; feng2024multiscale ; wei2024diffusion ; li2023hyperspectral . Some works also apply DMs to tasks like change detection bandara2022ddpm and climate prediction gao2023prediff , enabling better integration of multimodal data and improving pixel-level discrimination. On the other hand, image synthesis methods focus on generating artificial data to address the issue of data scarcity. Recent studies tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth demonstrate how synthetic data can benefit downstream remote sensing tasks. Unlike prior approaches that fine-tune on a single-source dataset for one specific task, we adopt a more general training and synthesis strategy. Our approach supports multiple tasks—scene classification, object detection, and semantic segmentation—allowing the model to generate more diverse and widely applicable data.
Figure 2: EarthSynth is trained with CF-Comp training strategy on real and unrealistic logical mixed data distribution, learns remote sensing pixel-level properties in multiple dimensions, and builds a unified process for conditional diffusion training and synthesis.
Generative Data Augmentation with Diffusion Models. Generative data augmentation using DMs has been explored across various domains, including natural images and remote sensing. Studies such as he2022synthetic ; trabucco2023effective investigate using text-to-image DMs to boost classification performance under limited data conditions. To further improve generation quality, alemohammad2024self ; alemohammad2024selfconsuming propose leveraging self-generated data; however, this approach risks model collapse due to overfitting. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. In remote sensing, recent works tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth have begun to explore the potential of diffusion-based data synthesis and enhancement to improve RSI interpretation tasks, like object detection or segmentation. But there is no unified solution to achieve multi-task generation.
@@ 3 Preliminaries
where $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$ consists of a conditional Mask-Text pairs $c_{i}$, and each conditional distribution $\mathbb{P}_{G_{\theta}}(x\mid c_{i})$ is induced from $x=G_{\theta}(\epsilon\mid c_{i})$ by sampling $\epsilon\sim\mathcal{N}(0,I)$. The pixel-level, region-level and semantic-level label $y=\{y_{l},y_{c}\}=\mathcal{F}(c_{i})$ is generated via a mapping function $\mathcal{F}$, where $y$ consists of a location label $y_{l}$ (semantic mask or bounding box) and a category label $y_{c}$.
@@
[…]
where $y$ depends on object $x_{\texttt{obj}}$ and background $x_{\texttt{bg}}$ but is independent of noise $x_{\texttt{noise}}$ which yields $\mathbb{P}_{G_{\theta}}(y\mid x)=\mathbb{P}_{G_{\theta}}(x_{\texttt{obj}},x_{\texttt{bg}})$. $x_{\texttt{noise}}$ is the noise disturbance during satellite imaging that makes semantic confusion pan2023reducing ; pan2023prior , leading to inter-class similarity and significant intra-class variation of satellite images.
[…]
@@ 4 Counterfactual Composition for Controllable Diffusion Training
Due to the inherent characteristics of satellite imagery, satellite images often exhibit high inter-class similarity and significant intra-class variation, posing challenges for RSI interpretation. To get more informative data distribution from a generative DM, we aim to approximate a real-world distribution $\mathcal{D}^{\texttt{real}}$ with as much training data distribution $\mathcal{D}^{\texttt{train }}\subset\mathcal{D}^{\texttt{real}}$ as possible. Constructing training data with diverse backgrounds and objects enables the DM to learn rich and diverse semantic information in open remote sensing scenarios. To achieve this, we enhance scene diversity through counterfactual composition, which involves combining existing object categories with diverse background contexts. The definition of counterfactual composition is as follows.
Figure 3: Copy-Paste operation of CF-Comp Strategy and Rule-based Filtering with CLIP.
@@
Counterfactual Composition allows for the generation of novel and logically consistent scenes by fusing diverse elements. The goal is to preserve semantic and structural coherence while expanding the diversity and improving generalization of possible inputs.
Unrealistic Logical Scene. We aim to construct an unrealistic logical scene by counterfactual composition, i.e., satellite imagery of real-world data that does not exist in the real world but is logically present. This also intended to maintain consistency in the image-label union space and prevent disruptions caused by arbitrary counterfactual composition. As shown in Figure 2, we use the Copy-Paste to perform counterfactual composition. For example, the unrealistic logical scene is obtained by combining the two images’ small and large vehicle objects in Figure 3. An unrealistic logical scene can be judged from three dimensions: channel, pixel, and semantic space. We define three criteria to measure whether two images can be combined into an unrealistic logical scene:

### v2

@@ Abstract
Remote sensing image (RSI) interpretation typically faces challenges due to the scarcity of labeled data, which limits the performance of RSI interpretation tasks. To tackle this challenge, we propose EarthSynth, a diffusion-based generative foundation model that enables synthesizing multi-category, cross-satellite labeled Earth observation for downstream RSI interpretation tasks. To the best of our knowledge, EarthSynth is the first to explore multi-task generation for remote sensing, tackling the challenge of limited generalization in task-oriented synthesis for RSI interpretation. EarthSynth, trained on the EarthSynth-180K dataset, employs the Counterfactual Composition training strategy with a three-dimensional batch-sample selection mechanism to improve training data diversity and enhance category control. Furthermore, a rule-based method of R-Filter is proposed to filter more informative synthetic data for downstream tasks. We evaluate our EarthSynth on scene classification, object detection, and semantic segmentation in open-world scenarios. There are significant improvements in open-vocabulary understanding tasks, offering a practical solution for advancing RSI interpretation.
@@ 1 Introduction
In parallel, recent advances in generative data augmentation antoniou2017data , particularly Diffusion Models (DMs) rombach2022high , offer a promising avenue for synthesizing high-quality labeled training data. Data augmentation using DMs has been widely applied across various domains chen2024comprehensive . Prior works he2022synthetic ; trabucco2023effective investigate text-to-image (T2I) models huang2024learning to boost classification performance in data-scarce scenarios, enlightening and inspiring the remote sensing community. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. By augmenting rare classes and enriching data diversity NEURIPS2023_f99f7b22 , these models can play a critical role in mitigating data scarcity and enhancing the generalization capabilities of RSI interpretation. Unlike existing studies, our work centers on improving the generative diversity of diffusion models tailored for remote sensing applications.
@@
In the remote sensing community, most existing data augmentation approaches are trained predominantly on single-source EO, which inherently limits the diversity of generated categories and lacks generality and flexibility across broader remote sensing applications. However, training generative models on multi-source data requires balancing the quality and diversity of generated data. Txt2Img-MHN xu2023txt2img attempts to use GANs salimans2016improved to generate satellite images. DiffusionSat khanna2023diffusionsat and CRS-Diff tang2024crs propose conditional DMs for generating optical RSIs, incorporating diverse texture-based conditions and applying the synthesized data to downstream tasks such as road extraction. GeoSynth toker2024satsynth joint data manifold of images and labels for satellite semantic segmentation. AeroGen tang2024aerogen and MMO-IG yang2025mmo try to use synthesized training data for satellite object detection. However, these methods are typically trained on single tasks or single-source data, requiring repeated training and generation to meet diverse needs, which hinders their application in real-world scenarios.
@@
To solve the above problems, we propose EarthSynth, a diffusion-based generative foundation model, synthesizing EO with a semantic mask and text for downstream RSI interpretation tasks as shown in Figure 1. First, we construct the EarthSynth-180K with multi-source and multi-category data, 180K samples for training EarthSynth. Specifically, we collect samples from multiple datasets and apply random cropping and category-augmentation strategy to standardize image resolution, ensuring alignment among images, semantic masks, and text descriptions. To our knowledge, EarthSynth-180K is the first large-scale remote sensing dataset for diffusion training. During training, we adopt the Counterfactual Composition (CF-Comp) strategy with channel, pixel, and semantic spaces as batch-sample selection mechanism to simultaneously enhance layout controllability and category diversity, thereby enabling the generation of more informative EO data. Different previous studies, we apply the CF-Comp strategy to diffusion-based generative foundation model training for downtown tasks, avoiding repeating the training task-specific generative model for each downstream task. For training data synthesis, a rule-based method of R-Filter is proposed to filter more informative synthetic data. We evaluate our EarthSynth with multiple datasets on scene classification, object detection, and semantic segmentation. Furthermore, the effectiveness of our method is demonstrated through comprehensive ablation studies and visualization analysis.
@@
  • We propose EarthSynth, a diffusion-based generative foundation model trained on the EarthSynth-180K dataset with 180K cross-satellite and multi-sensor samples aligned image, semantic mask, and text, achieving a unified solution to achieve multi-task generation.
  • EarthSynth employs the CF-Comp strategy to balance the layout controllability and category diversity during training, enabling fine layout control for RSI generation. And integrates the R-Filter post-processing method to extract more informative synthesized data.
  • EarthSynth is evaluated on scene classification, object detection, and semantic segmentation in open-vocabulary scenarios, well validating its effectiveness.
@@ 2 Related Work
Diffusion Models for Remote Sensing. Diffusion Models (DMs), which have shown great success in natural image synthesis rombach2022high , are increasingly being applied to remote sensing khanna2023diffusionsat ; liu2025text2earth . Depending on the use case liu2024diffusion , their applications in remote sensing can be broadly divided into three categories: image enhancement, image interpretation, and image synthesis. For image enhancement, DMs improve multispectral and hyperspectral images by fusing information across channels and expanding attention xu2023dual ; feng2024multiscale ; wei2024diffusion ; li2023hyperspectral . Some works also apply DMs to tasks like change detection bandara2022ddpm and climate prediction gao2023prediff , enabling better integration of multimodal data and improving pixel-level discrimination. On the other hand, image synthesis methods focus on generating artificial data to address the issue of data scarcity. Recent studies tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth demonstrate how synthetic data can benefit downstream remote sensing tasks. Unlike prior approaches that fine-tune on a single-source dataset for one specific task, we adopt a more general training and synthesis strategy. Notably, these data augmentation methods zhang2024advancing ; zhang2025controllable ; liu2025control ; li2025domain demonstrate significant advantages in few-shot learning tasks fu2023styleadv ; fu2025ntire . Our approach supports multiple tasks, including scene classification, object detection, and semantic segmentation, allowing the model to generate more diverse and widely applicable data.
Figure 2: EarthSynth is trained with CF-Comp training strategy on real and unrealistic data distribution, learns remote sensing pixel-level properties in multiple dimensions, and builds a unified process for conditional diffusion training and synthesis.
Generative Data Augmentation with Diffusion Models. Generative data augmentation using DMs has been explored across various domains, including natural images and remote sensing. Studies such as he2022synthetic ; trabucco2023effective investigate using text-to-image DMs to boost classification performance under limited data conditions. To further improve generation quality, alemohammad2024self ; alemohammad2024selfconsuming propose leveraging self-generated data; however, this approach risks model collapse due to overfitting. And some efforts alemohammad2024self ; alemohammad2024selfconsuming research the risk of model collapse and improve generation quality through self-generated data. In remote sensing, recent works tang2024crs ; toker2024satsynth ; tang2024aerogen ; yang2025mmo ; sastry2024geosynth ; li2025domain have begun to explore the potential of diffusion-based data synthesis and enhancement to improve RSI interpretation tasks. But there is no unified solution to achieve multi-task generation.
@@ 3 Preliminaries
where $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$ consists of a conditional Mask-Text pairs $c_{i}$, and each conditional distribution $\mathbb{P}_{G_{\theta}}(x\mid c_{i})$ is defined by generating samples $x=G_{\theta}(\epsilon\mid c_{i})$, where $\epsilon\sim\mathcal{N}(0,I)$. The pixel-level, region-level, and semantic-level label $y=\{y_{l},y_{c}\}=\mathcal{F}(c_{i})$ is generated via a mapping function $\mathcal{F}$, where $y$ consists of a location label $y_{l}$ (semantic mask or bounding box) and a category label $y_{c}$.
@@
$x_{\texttt{obj}},x_{\texttt{bg}}\not\!\perp\!\!\!\perp y,\quad x_{\texttt{noise}}\perp\!\!\!\perp y,\\$ (2)
where $y$ depends on object $x_{\texttt{obj}}$ and background $x_{\texttt{bg}}$ but is independent of noise $x_{\texttt{noise}}$ which yields $\mathbb{P}_{G_{\theta}}(y\mid x)=\mathbb{P}_{G_{\theta}}(x_{\texttt{obj}},x_{\texttt{bg}})$. $x_{\texttt{noise}}$ is the noise disturbance during satellite imaging that makes semantic confusion pan2023reducing ; pan2023prior ; pan2024pir , leading to inter-class similarity and significant intra-class variation of satellite images.
Copy-Paste Augmentation. Copy-Paste ghiasi2021simple is a data augmentation technique that involves copying objects or regions from one image and pasting them into another to create new composite scenes. As shown in Figure 3, $\texttt{Copy-Paste}(x^{a}_{obj},x^{b})$ represents the operation of copying the objects of image $x^{a}$ to image $x^{b}$. However, the Copy-Paste introduces compositional artifacts or non-smooth transitions, etc., that alter the statistical properties of the image distribution. These artifacts and transitions can typically be mitigated through the training process of DMs. More details are described in the Appendix.
@@ 4 Counterfactual Composition for Controllable Diffusion Training
Due to the inherent characteristics of satellite imagery, satellite images often exhibit high inter-class similarity and significant intra-class variation, posing challenges for RSI interpretation. To get more informative data distribution from a generative DM, we aim to approximate a real-world distribution $\mathcal{D}^{\texttt{real}}$ with as much training data distribution $\mathcal{D}^{\texttt{train}}\subset\mathcal{D}^{\texttt{real}}$ as possible. Constructing training data with diverse backgrounds and objects enables the DM to learn rich and diverse semantic information in open remote sensing scenarios. To achieve this, we enhance scene diversity through counterfactual composition, which involves combining existing object categories with diverse background contexts. The definition is as follows.
Figure 3: Left: Copy-Paste used in CF-Comp Strategy. Right: CLIP-based rule filtering retains high-quality images.
@@
Counterfactual Composition allows for the generation of novel and logically consistent scenes by fusing diverse elements. The goal is to preserve semantic and structural coherence while expanding the diversity and improving the generalization of possible inputs.